OpenStudy (anonymous):

if you want to find the velocity function of a problem would you just take the 1st derivative of the original equation? for ex. s(t)=-16t^2-64t+512?

OpenStudy (amistre64):

velocity is the derivative of position; yes the change of distance with respect to time/parameter

OpenStudy (anonymous):

yeah ds/dt = v

OpenStudy (anonymous):

ok now my next question is when you are asked to find the acceleration of an object how would you do that? after finding the 1st derivative/velocity function?

OpenStudy (anonymous):

differejntiate it again

OpenStudy (anonymous):

a=dv/dt so a = d2v/dv2

OpenStudy (anonymous):

so i should take the derivative of the velocity function?

OpenStudy (amistre64):

what is curvature :)

OpenStudy (amistre64):

the change in direction with respect to distance :)

OpenStudy (anonymous):

no its the beauty of jennifer aniston..lol

OpenStudy (anonymous):

yeah a=dv/dt

OpenStudy (amistre64):

does your s(t) there indicate position of speed?

OpenStudy (anonymous):

position as a function of time

OpenStudy (amistre64):

the derivative of that function will tell you the speed it is going at any given position in time then

OpenStudy (anonymous):

yea the reason why im asking this is because my professor gave us a question like this a while back and he asked for the velocity function, and then the acceleration. i got the velocity function no problem but then messed up and got the acceleration wrong. so just to clarify (velocity function = 1st derivative of original function, acceleration function = derivative of the velocity function?) thanks for all the help i really appreciate it

OpenStudy (anonymous):

yeah its right

OpenStudy (amistre64):

the derivative of velocity as part of acceleration.... if you do vectors

OpenStudy (amistre64):

a = d^2s/dt^2.<T> + k(ds/dt)^2<N>

OpenStudy (anonymous):

ok thanks everyone, im sure i will have a few more questions as this study session progressively gets more intense!